# What is the Best Way to Do Something? A Discreet Tour of Discrete Optimization

### v1

@@
Abstract
In mathematical optimization, we want to find the best possible solution for a decision-making problem. Curiously, these problems are harder to solve if they have discrete decisions. Imagine that you would like to buy chocolate: you can buy no chocolate or one chocolate bar, but typically you cannot buy just half of a bar. Now imagine that you could also buy many other items, and that you need to meet nutritional needs while minimizing the grocery bill. With more options and more demands, finding the best solution becomes trickier. But since many real-world settings benefit from mathematical optimization, such as scheduling trains and flights, planning truck deliveries, and making better investment decisions, these problems are widely studied in a branch of mathematics called Operations Research (OR). Sometimes we can simply write the mathematical model and find an optimal solution with OR software, but for larger problems we may need to develop new mathematical models and even write our own algorithms. We explore both cases with a simple and well-known problem (the traveling salesperson problem), some computer programming (in Python), and software that is free for academic use (Gurobi).
@@ Abstract
All the code and data used will be available at: https://github.com/thserra/discreet
@@ 1 Introduction
Second, if we do not have to buy whole fruit, then the problem becomes finding the cheapest source of potassium per dollar, regardless of how much we need. With apples, you get $\frac{214}{0.49}\approx 398$ mg of potassium per dollar. With oranges, you get $\frac{232}{0.50}\approx 458$ mg of potassium per dollar. Hence, buying $\frac{235}{232}\approx 1.013$ oranges (for $\approx\$0.51$) would be the cheaper than buying $\frac{235}{214}\approx 1.010$ apples (for $\approx\$0.54$).
@@
In the example above, the optimal solution with integer variables is to buy two apples for \$0.98\\$0.98. However, the optimal solution with continuous variables is to buy $\approx 1.013$ oranges for $\approx\$0.51$, which is much cheaper. We call the value of the objective function in an optimal solution as the optimal value. If the only difference between two minimization models $A$ and $B$ is that the decision variables are continuous in $A$ and integer in $B$, is the optimal value for $A$ always smaller?
@@
    Now suppose that we only get paid \$0.48\\$0.48 for our orange, even though we have to pay \$0.50\\$0.50 to buy another one. This is what happens when you trade currency, say US dollars for Brazilian reais: the shop buys for less than they sell, since maintaining a shop costs money and they also need to profit for a living. Digression apart, how can you model having an orange to sell along with buying apples and oranges, such that we spend the least to meet the potassium need?
@@ 3 The Branch-and-Bound Algorithm
To describe how branch-and-bound works, we will use it to solve the four fruit model from Section 2. If we solve the LP relaxation of that model, we find an optimal solution $(x_{1},x_{2},x_{3},x_{4})\approx(0.00,0.95,0.00,0.53)$ with the optimal value $z\approx 0.75$. That is not a feasible solution for the ILP model. However, it shows that no solution has a better value than \$0.75\\$0.75. This is a lower bound on the optimal value.
@@
By solving the models of the children nodes, we keep finding noninteger values and still have no feasible solution. However, we find out that there is no feasible solution better than \$0.78\\$0.78, which is the optimal value of the LP relaxation of the left child, since the optimal value of the LP relaxation of the right child is greater than that. Hence, we can update the lower bound to $0.78$, which is more accurate.
@@
This time we are luckier. From the third explored node in Figure 3, the left child (a rectangle) has a feasible solution of buying two oranges for \$1.00\\$1.00. Hence, we now have an upper bound on the optimal value. The right child of node 3 happens to produce a new lower bound of \$0.93\\$0.93, since it becomes the smallest optimal value of the LP relaxation among all unexplored nodes. Moreover, we can now discard the right child of node 2, since we would not be able to find a better solution from it than the one that we already have. We say that this node has been fathomed.
The bounding step, through which we reduce the gap between the upper and the lower bounds, is what prevents us from doing an almost [^10] We will return to this point in Exercise 7. complete enumeration of feasible solutions with branch-and-bound. Still, we have \$0.07\\$0.07 left between the best solution found and the theoretical best solution at this point, so we must proceed!
@@
We branch on the right child of node 3 in Figure 4, through which we obtain a left node that is automatically fathomed due to $z$ and a right node producing a better feasible solution of buying one apple and one orange for \$0.99\\$0.99, hence lowering the upper bound. As we have explored the node with a lower bound of \$0.93\\$0.93, we raise the lower bound to the value of the only unexplored node at this point: \$0.94\\$0.94.
Now we have only one unexplored node (the right child of node 1) with only one noninteger variable to branch with ($x_{2}$), which when explored becomes node 5 in Figure 5. We obtain another feasible solution from it, but both child nodes are fathomed due their values. Since we do not have unexplored nodes left, we conclude that the optimal value is \$0.99\\$0.99, with $(x_{1},x_{2},x_{3},x_{4})=(1,1,0,0)$ as an optimal solution.
@@
    Buying two apples as in an optimal solution of our first model in Section 1 would have been cheaper than \$0.99\\$0.99. Why can’t we do that now?
@@ 5 Assignment Problems and Graphs
  • In Challenge Problem 2, we had a facility location problem in which we had to decide which facilities to rent and, based on those, which facility would supply each store. The decision variables were provided to make it easier to understand how to model it. Still, this was a challenge problem because reasoning about constraints involving binary variables is different than thinking about quantities.202020 Going back to the last item in part (a) of Challenge Problem 2, you can use the constraint $x_{ij}\leq y_{i}$ to ensure that $x_{ij}=1$ only if $y_{i}=1$: If $y_{i}=0$, then $x_{ij}=0$. If $y_{i}=1$, then $x_{ij}$ is not constrained.
@@ 5.1 The First Model: One-Way Assignment
To start modeling the problem, let us consider the fact that every employee needs to travel to another office. We start with three offices, numbered from 1 to 3, and we use $d_{ij}$ as the distance from $i$ to jj.212121 If $d_{ij}\neq d_{ji}$, we say that the problem is assymetric. However, our example will be symmetric. That can be modeled as follows:
@@ 7 Modeling Tricks and Model Strength
In practice, we say that one model is stronger than another for the same problem if it has fewer fractional solutions. As seen in Chapter 3, branch-and-bound can be a lengthy process if too many LP relaxations have solutions with fractional values for the integer variables of the ILP model, to the point of taking an exponential time on the size of the instance. Note that using canonical cuts on the unit hypercut is a first step, but it may not be sufficient: It is possible to find a constraint that removes even more fractional solutions if one of the neighbors of $\bar{\textbf{x}}$ is infeasible. More generally, we need to compare entire ILP models for the same problem, in the hope of finding which ones have the smallest LP relaxations.292929 Whereas solutions on integer variables are countable and finite if the decision variables are bounded, solutions on continuous variables are not. Let’s think of those as a continuous set. To focus on models that we can compare directly, consider the definitions below:
@@
Since $\textbf{x}\geq 0$, then the subtour elimination constraint implies both cuts used for eliminating specific subtours. On the other hand, neither of the cuts removing subtours would remove $(x_{12},x_{13},x_{21},x_{23},x_{31},x_{32})=(0.5,0.5,0.5,0.5,0.5,0.5)$, which is a fractional solution removed by the classic subtour elimination constraint. Therefore, the TSP Model with the classic subtour elimination constraints strictly dominates an alternate model with cuts removing each subtour. Better yet, the stronger model also has fewer constraints!303030 Although model strength is important, an excessive number of constraints may certainly slow down the time for solving the LP relaxation at each node of the branch-and-bound tree. According to Bill Cook, sometimes solving the LP relaxation is the bottleneck in large TSP models. That may also be an issue in cases where we use constraints to reduce the feasible set when there is symmetry in the model. These are called symmetry-breaking constraints Margot2010 and they often help solving the problem faster, but in some cases the number of such constraints is so large that adding them to the model increases the runtime. One alternative in those cases is to only exploit the symmetries by building a different type of branch-and-bound tree, such as with orbital branching orbital .
@@ 9 Designing Your Own (Heuristic) Algorithms
When running this code for the whole 1000 cities, you may feel a deep satisfaction in how fast this turns out to be. But, of course, everything comes with a price: not every choice for this tour will seem to be a reasonable one afterwards. Back in the 1970s and 1980s, when finding optimal solutions was very difficult even for reasonably sized problems, many interesting ideas came up about what to do instead. In fact, these ideas remain alive and new ones are still being proposed because we can always find a larger or more complicated problem to solve.323232 We discuss alternative approaches with historical references from when they were proposed as well as more recent references on their application. You may need them one day. Besides, this section is more generours in citations than the previous ones because we cover other approaches here. For more references on the topics covered before in the text, see the next (and last) section.
@@ 9.3 Quantum Computing and Machine Learning
See the instructions in other challenge problems and their footnotes for guidance on how to conduct similiar experiments and implement randomization in Python. You may also consider using scatter plots as in Exercise 27, provided that the comparisons presented in such plots are reasonable.353535 This is a situation in which we risk “comparing apples and oranges” if we are not careful, in the sense that the results are not meaningful. Comparing the runtime of solving an ILP model to optimality with the time to generate a solution that is not optimal would not make sense. However, you may consider comparing the former with the time by which the first optimal solution is generated with an heuristic (even if we cannot tell if a solution is optimal by using the heuristic alone, in this case we would know that from having solved the ILP model and that helps us evaluate how good the heuristic is). From a computational perspective, that would tell us if putting the same amount of computational effort would possibly produce an optimal solution with an heuristic.

### v3

@@
11institutetext: Thiago Serra 22institutetext: University of Iowa, United States, 22email: thiago-serra@uiowa.edu
What is the Best Way to Do Something?
A Discreet Tour of Discrete Optimization
Thiago Serra
@@ A Discreet Tour of Discrete Optimization
All the code and data used is available at: https://github.com/thserra/discreet
@@ 1 Introduction
Second, if we do not have to buy whole fruit, then the problem becomes finding the cheapest source of potassium per dollar, regardless of how much we need. With apples, you get $\frac{214}{0.49}\approx 398$ mg of potassium per dollar. With oranges, you get $\frac{232}{0.50}\approx 458$ mg of potassium per dollar. Hence, buying $\frac{235}{232}\approx 1.013$ oranges (for $\approx\mathdollar 0.51$) would be the cheaper than buying $\frac{235}{214}\approx 1.010$ apples (for $\approx\mathdollar 0.54$).
@@
In the example above, the optimal solution with integer variables is to buy two apples for $\mathdollar 0.98$. However, the optimal solution with continuous variables is to buy $\approx 1.013$ oranges for $\approx\mathdollar 0.51$, which is much cheaper. We call the value of the objective function in an optimal solution as the optimal value. If the only difference between two minimization models $A$ and $B$ is that the decision variables are continuous in $A$ and integer in $B$, is the optimal value for $A$ always smaller?
@@
    Now suppose that we only get paid $\mathdollar 0.48$ for our orange, even though we have to pay $\mathdollar 0.50$ to buy another one. This is what happens when you trade currency, say US dollars for Brazilian reais: the shop buys for less than they sell, since maintaining a shop costs money and they also need to profit for a living. Digression apart, how can you model having an orange to sell along with buying apples and oranges, such that we spend the least to meet the potassium need?
@@ 3 The Branch-and-Bound Algorithm
To describe how branch-and-bound works, we will use it to solve the four fruit model from Section 2. If we solve the LP relaxation of that model, we find an optimal solution $(x_{1},x_{2},x_{3},x_{4})\approx(0.00,0.95,0.00,0.53)$ with the optimal value $z\approx 0.75$. That is not a feasible solution for the ILP model. However, it shows that no solution has a better value than $\mathdollar 0.75$. This is a lower bound on the optimal value.
@@
By solving the models of the children nodes, we keep finding noninteger values and still have no feasible solution. However, we find out that there is no feasible solution better than $\mathdollar 0.78$, which is the optimal value of the LP relaxation of the left child, since the optimal value of the LP relaxation of the right child is greater than that. Hence, we can update the lower bound to $0.78$, which is more accurate.
@@
This time we are luckier. From the third explored node in Figure 3, the left child (a rectangle) has a feasible solution of buying two oranges for $\mathdollar 1.00$. Hence, we now have an upper bound on the optimal value. The right child of node 3 happens to produce a new lower bound of $\mathdollar 0.93$, since it becomes the smallest optimal value of the LP relaxation among all unexplored nodes. Moreover, we can now discard the right child of node 2, since we would not be able to find a better solution from it than the one that we already have. We say that this node has been fathomed.
The bounding step, through which we reduce the gap between the upper and the lower bounds, is what prevents us from doing an almost [^10] We will return to this point in Exercise 7. complete enumeration of feasible solutions with branch-and-bound. Still, we have $\mathdollar 0.07$ left between the best solution found and the theoretical best solution at this point, so we must proceed!
@@
We branch on the right child of node 3 in Figure 4, through which we obtain a left node that is automatically fathomed due to $z$ and a right node producing a better feasible solution of buying one apple and one orange for $\mathdollar 0.99$, hence lowering the upper bound. As we have explored the node with a lower bound of $\mathdollar 0.93$, we raise the lower bound to the value of the only unexplored node at this point: $\mathdollar 0.94$.
Now we have only one unexplored node (the right child of node 1) with only one noninteger variable to branch with ($x_{2}$), which when explored becomes node 5 in Figure 5. We obtain another feasible solution from it, but both child nodes are fathomed due their values. Since we do not have unexplored nodes left, we conclude that the optimal value is $\mathdollar 0.99$, with $(x_{1},x_{2},x_{3},x_{4})=(1,1,0,0)$ as an optimal solution.
@@
    Buying two apples as in an optimal solution of our first model in Section 1 would have been cheaper than $\mathdollar 0.99$. Why can’t we do that now?
@@ 5 Assignment Problems and Graphs
  • In Challenge Problem 2, we had a facility location problem in which we had to decide which facilities to rent and, based on those, which facility would supply each store. The decision variables were provided to make it easier to understand how to model it. Still, this was a challenge problem because reasoning about constraints involving binary variables is different than thinking about quantities. [^20] Going back to the last item in part (a) of Challenge Problem 2, you can use the constraint $x_{ij}\leq y_{i}$ to ensure that $x_{ij}=1$ only if $y_{i}=1$: If $y_{i}=0$, then $x_{ij}=0$. If $y_{i}=1$, then $x_{ij}$ is not constrained.
@@ 5.1 The First Model: One-Way Assignment
To start modeling the problem, let us consider the fact that every employee needs to travel to another office. We start with three offices, numbered from 1 to 3, and we use $d_{ij}$ as the distance from $i$ to jj. [^21] If $d_{ij}\neq d_{ji}$, we say that the problem is assymetric. However, our example will be symmetric. That can be modeled as follows:
@@ 7 Modeling Tricks and Model Strength
In practice, we say that one model is stronger than another for the same problem if it has fewer fractional solutions. As seen in Chapter 3, branch-and-bound can be a lengthy process if too many LP relaxations have solutions with fractional values for the integer variables of the ILP model, to the point of taking an exponential time on the size of the instance. Note that using canonical cuts on the unit hypercut is a first step, but it may not be sufficient: It is possible to find a constraint that removes even more fractional solutions if one of the neighbors of $\bar{\textbf{x}}$ is infeasible. More generally, we need to compare entire ILP models for the same problem, in the hope of finding which ones have the smallest LP relaxations. [^29] Whereas solutions on integer variables are countable and finite if the decision variables are bounded, solutions on continuous variables are not. Let’s think of those as a continuous set. To focus on models that we can compare directly, consider the definitions below:
@@
Since $\textbf{x}\geq 0$, then the subtour elimination constraint implies both cuts used for eliminating specific subtours. On the other hand, neither of the cuts removing subtours would remove $(x_{12},x_{13},x_{21},x_{23},x_{31},x_{32})=(0.5,0.5,0.5,0.5,0.5,0.5)$, which is a fractional solution removed by the classic subtour elimination constraint. Therefore, the TSP Model with the classic subtour elimination constraints strictly dominates an alternate model with cuts removing each subtour. Better yet, the stronger model also has fewer constraints! [^30] Although model strength is important, an excessive number of constraints may certainly slow down the time for solving the LP relaxation at each node of the branch-and-bound tree. According to Bill Cook, sometimes solving the LP relaxation is the bottleneck in large TSP models. That may also be an issue in cases where we use constraints to reduce the feasible set when there is symmetry in the model. These are called symmetry-breaking constraints Margot2010 and they often help solving the problem faster, but in some cases the number of such constraints is so large that adding them to the model increases the runtime. One alternative in those cases is to only exploit the symmetries by building a different type of branch-and-bound tree, such as with orbital branching orbital.
@@ 9 Designing Your Own (Heuristic) Algorithms
When running this code for the whole 1000 cities, you may feel a deep satisfaction in how fast this turns out to be. But, of course, everything comes with a price: not every choice for this tour will seem to be a reasonable one afterwards. Back in the 1970s and 1980s, when finding optimal solutions was very difficult even for reasonably sized problems, many interesting ideas came up about what to do instead. In fact, these ideas remain alive and new ones are still being proposed because we can always find a larger or more complicated problem to solve. [^32] We discuss alternative approaches with historical references from when they were proposed as well as more recent references on their application. You may need them one day. Besides, this section is more generours in citations than the previous ones because we cover other approaches here. For more references on the topics covered before in the text, see the next (and last) section.
@@ 9.3 Quantum Computing and Machine Learning
See the instructions in other challenge problems and their footnotes for guidance on how to conduct similiar experiments and implement randomization in Python. You may also consider using scatter plots as in Exercise 27, provided that the comparisons presented in such plots are reasonable. [^35] This is a situation in which we risk “comparing apples and oranges” if we are not careful, in the sense that the results are not meaningful. Comparing the runtime of solving an ILP model to optimality with the time to generate a solution that is not optimal would not make sense. However, you may consider comparing the former with the time by which the first optimal solution is generated with an heuristic (even if we cannot tell if a solution is optimal by using the heuristic alone, in this case we would know that from having solved the ILP model and that helps us evaluate how good the heuristic is). From a computational perspective, that would tell us if putting the same amount of computational effort would possibly produce an optimal solution with an heuristic.